OpenStudy (anonymous):

So. How much will 5 quarters and 6 dimes cost if I add them together? :/

OpenStudy (shadowlegendx):

Quarter = .25 Dime = .10 5(.25) + 6(.10) = Your Answer

OpenStudy (shadowlegendx):

Can you solve that equation?

OpenStudy (shadowlegendx):

If you can't, that's completely fine and I can help you further :)

OpenStudy (anonymous):

Yes, please help. :)

OpenStudy (shadowlegendx):

\[5 \times .25 \] or \[5(.25) \] \[= 1.25\]

OpenStudy (shadowlegendx):



OpenStudy (shadowlegendx):

\[6 \times .10\] or \[6(.10)\] \[=.60\]| This is pretty simple as these are just base numbers

OpenStudy (shadowlegendx):

So now we have \[1.25 + .60 = ?\] Can you solve that?

OpenStudy (mathmale):

How much will 5 quarters and 6 dimes cost if I add them together? would be much clearer as: What is the total value of 5 quarters and 6 dimes? A coin doesn't "cost" some figure; instead, a coin has a "value."

OpenStudy (anonymous):

Yes, ok thank you'! ^^ and sorry I'm in 5th grade so yeah'.... I can take it from here . Thank you very much :) Shadowlegendx

OpenStudy (shadowlegendx):

1.85 should be your answer :)

OpenStudy (mathmale):

Total value = value of 1 dime times the number of dimes PLUS value of 1 quarter times the number of quarters.

OpenStudy (mathmale):

Use the dollar sign here.

OpenStudy (shadowlegendx):

$1.85 :P

OpenStudy (shadowlegendx):

Thank you for pointing that out @mathmale Always use the dollar sign in your answer when it's money :P